OpenStudy (anonymous):

Identify the CENTER and the RADIUS of the circle described by the equation ... 2x^2 + 2y^2 - 12x - 16y = 0

OpenStudy (anonymous):

\[2x^2 + 2y^2 - 12x - 16y = 0\] to make it look like \[(x-h)^2+(y-k)^2=r^2\] is an exercise in completing the square twice